I just really dont understand math at all im having trouble in my algebra2 class with point slope heres a problem can u solve it for me [3,2] [-4,3]
Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

This site gives a good tutorial on the subject:
http://www.purplemath.com/modules/slope.htm
.
Basically, the formula is:
slope = (y2-y1)/(x2-x1)
where
the x's and y's are the given points.
.
For example, your two points:
[3,2] [-4,3]
is mapped to:
[x1,y1] [x2,y2]
.
Then
slope = (y2-y1)/(x2-x1)
would be:
slope = (3-2)/(-4-3)
slope = (1)/(-7) = -1/7
.
That's it... your slope between those two points is -1/7

Answer by Alan3354(69443)

I just really dont understand math at all im having trouble in my algebra2 class with point slope heres a problem can u solve it for me [3,2] [-4,3]
--------------------
You didn't say what you want to do. "Solve it" is a general term.
I'll take a guess you want to find the equation of the line that goes thru the 2 points, and express it in slope-intercept form. But, a good first step in any problem, math, personal, or otherwise, is to define what you want to do.
----------------
Problem: Find the eqn of the line that passes thru the points [-4,3] and [3,2]
Use the slope-intercept eqn to start, and find the slope:
m = (y2-y1)/(x2-x1)
m = (2-3)/(3 - (-4))
m = (-1)/(7)
m = -1/7 (the slope of the line)
Now use either point to find the eqn thru the 2 points.
y-y1 = m*(x-x1)
y-3 = (-1/7)*(x-(-4))
y-3 = (-1/7)*(x+4)
y-3 = (-1/7)x - 4/7
y = (-1/7)x + 17/7 Slope-intercept form, y = mx + b
Rearrange to "standard form"
7y = -x + 17
x + 7y = 17
----------
If you check, you'll see that both points satisfy the eqn.